# Educating LLMs like Human Students: Structure-aware Injection of Domain Knowledge

### v1

@@ 4 Experiments
We design a comprehensive evaluation of our StructTuning through seveal experiments on two benchmarks. First, we investigate the free-form question-answering task on the LongBench [4] dataset, in order to verify the memorization and understanding of injected knowledge (the answer can be directly found in training corpora). Then, we delve into the multi-choice question-answering task on MMedBench [35], to explore how LLMs apply the injected knowledge in basic medicine to determine the real-world diagnosis for patients with logical reasoning.
@@ 4.1 Preliminary Investigation on Free-form Question-Answering
Datasets and Tasks. LongBench [4] is a multi-task benchmark tailored for open-book reading comprehension evaluation, where LLMs generate answers to given questions based on one or several lengthy passages. To focus on knowledge injection, we turn the open-book evaluation into a closed-book QA task, where contextual passages are injected into LLMs with training techniques, and models answer questions without input passages. We choose 7 subsets with 1,350 test examples from LongBench for single- and multi-document question-answering evaluation, and the remaining synthetic or code-orientated tasks are eliminated. More details are described in Sec. A.1.
Evaluation Metrics. For free-form QA in LongBench [4], we report the F1-Score between models’ outputs and ground-truth answers for both open-book and closed-book QA tasks. Besides, we also report the Recall for model responses to quantify the memorization degree of injected knowledge.
Investigated Models. For LongBench, we mainly investigate the knowledge injection to the Llama2-7B model [40], and report the Open- and Closed-Book QA performance for comparison.
Implementation Details. On LongBench for closed-book QA, we continually pre-train the Llama2-7B model on 10,476 passages for 3 epochs using a batch size of 128. The initial learning rate is 2e-5, decayed to 0 at the end of training with a cosine scheduler. Then, we query Llama3-70B to generate 2,700 SSFT QA examples, and eliminate the generated data whose answer shares over 0.5 F1-Score with one of the test samples to avoid data leakage. We train the knowledge-injected models for 1 epoch to avoid overfitting. Models are trained with the LLaMA-Factory framework  [^2] https://github.com/hiyouga/LLaMA-Factory.
Figure 6: An example for structure-aware responses.
Main Results. In Tab. 1, we first report the Open-Book QA (OBQA) baseline for Llama2-7B with passages inputs, which shows the best performance on MultiFieldQA (MFQA) [4] and 2WikiMultihopQA (2Wiki) [17] subsets. Then, we establish the Closed-Book QA (CBQA) baseline for pre-trained Llama2-7B, which aims to identify the knowledge that has already been equipped from the newly injected knowledge. Consequently, we follow the traditional CPT+SFT paradigm to inject passage contents into model parameters, and supplement the experiment of our SCPT+SSFT technique for comparison. According to the results shown in Tab. 1, the traditional continual pre-training strategy immediately improves the QA performance on six out of seven subsets (except the Qasper [10] subset), while the supervised fine-tuning alignment does not bring significant gains on the overall F1-Score measure. In contrast, our SCPT+SSFT approach successfully boosts the closed-book QA performance to 20.1% on average, even surpassing the open-book QA baseline of 18.7%. It indicates the vanilla SFT strategy can only regularize LLMs’ response styles, while our SSFT could teach LLMs to utilize their knowledge (injected in the SCPT stage) to answer corresponding questions, as exemplified in Fig. 6.
Table 1: F1 Score evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench [4] dataset. The best results are marked in bold, and the secondary results are marked with underlines. The backbone model is Llama2-7B [40].
@@
OBQA - 19.2 36.8 11.9 25.4 32.8 9.4 5.2 18.7
CBQA - 17.3 15.9 1.4 5.7 6.7 3.3 1.8 7.4
CPT 16.8 23.6 17.5 18.0 18.6 6.7 14.3 16.5
CPT+SFT 16.8 23.1 13.2 21.3 19.1 10.4 13.4 16.8
SCPT 15.2 21.5 15.2 14.9 19.8 5.6 14.1 15.2
SCPT+SSFT 19.7 23.5 19.5 26.4 24.1 12.2 15.4 20.1
Table 2: Recall evaluation of the Closed-Book QA task. The best results are marked in bold.
Adaptation SingleDoc-QA MultiDoc-QA Average
Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
CPT 13.4 33.0 16.6 28.1 27.7 15.9 13.1 21.1
CPT+SFT 20.7 35.3 20.6 29.9 32.1 18.9 12.0 24.2
SCPT 18.8 42.5 17.7 35.7 36.4 20.5 15.3 26.7
@@
To further support this claim, we compute Recall between model outputs and ground-truth answers in Tab. 2 as a proxy measure for memorizing and manipulating the injected knowledge [59]. Specifically, we use a CoT instruction of Answer the question and explain why. to elicit models’ knowledge in their responses. As shown in Tab. 2, our SCPT strategy has already achieved higher knowledge recall than the CPT-SFT paradigm (26.7% v.s. 24.2%). It implies our structure-aware continually pre-trained model has successfully associated the relevant passages with their entire knowledge structure for the given question, which provides the knowledge path to seek targeted information to derive the answer. As displayed in Fig. 6, our SSFT technique explicitly teaches the model to recall the learned knowledge and answer the questions, which further improves the knowledge recall in Tab. 2.
@@
In addition, we also statistically quantify the memorization of injected knowledge structures by comparing the mindmap in models’ responses with the ground-truth answers via lexical ROUGE-L [27] and semantic BERTScore [54] measures. The results shown in Fig. 7 indicate a relatively good memorization of the injected knowledge mindmap, showing the efficacy of our SCPT strategy.
@@ 4.2 In-depth Evaluation for Multi-choice QA Application
Datasets and Tasks. MMedBench [35] is a multilingual medical multi-choice QA benchmark, with 45,048 QA pairs for adapting LLMs to medical experts and 8,518 for testing. We collect 76M textbook corpora from MedTextBooks [21] and MMedC [35] for medical knowledge injection and use the training/test split from MMedBench for SFT/evaluation. Detailed setup is in Sec. A.2.
@@
Evaluation Metrics. For the multi-choice QA task in MMedBench [35], we follow the default setting to calculate the accuracy on six language subsets, as well as the averaged scores.
@@
Investigated Models. For MMedBench, we extend the investigated LLMs across model scales and architectures including Llama2-7B, Llama2-13B [40], and InternLM2-7B [57]. We also compare our knowledge-injected models with other open- or closed-source LLMs, such as the domain-specified MedAlpaca [15], ChatDoctor [53], PMC-LLaMA [47], and MMedLM [35] models.
@@
Implementation Details. We first train LLMs for 3 epochs on medical textbooks with a batch size of 128. Then, we ask Llama3-70B to create structure-aware explanations for existing 45K QA samples in MMedBench’s training split and 33K extra entries by traversing extracted knowledge structures. Syntheses with overlapped options in the test set are removed. In the SFT phase, the learning rate is set as 1e-6 to avoid overfitting on such an amount of SFT samples, as suggested by Qiu et al. [35].
@@
Main Results. In Tab. 3, we present the overall performance across a series of LLMs on the testing split of MMedBench. The results demonstrate the promising enhancement achieved by our StructTuning technique, which translates into consistently significant improvements on Llama2-7B (+8.78%), Llama2-13B (+6.17%), and InternLM2-7B (+4.46%). It shows the generalizability and scalability of our StructTuning strategy across model architectures and sizes. Notably, our method, comprising SCPT on 76M curated textbook corpora and SSFT on 78K QA samples, achieves 50% performance (4.46% v.s. 8.76%) against the state-of-the-art MMedLM2-7B model, which is trained on a huge MMedC [35] corpora of 25.5B tokens. Our approach shows much more efficiency in transforming pre-trained LLMs into domain experts by structured-aware knowledge injection, which largely reduces the cost for LLMs’ domain adaptation to 0.3%. However, our method is currently unable to achieve 100% of knowledge injection efficacy against MMedLM2, which may comprise two major reasons: (1) some of the knowledge in the training corpus has already been learned by the foundation InternLM2 model, and (2) the knowledge in our curated 76M training data cannot cover all the testing samples. We leave the promotion of achieving 100% improvement in our future work.
@@
Table 3: Multiple-choice accuracy evaluation on MMedBench [35]. We report each model’s accuracy across six languages separately, with “Average” denoting the mean score over six languages. We also list out the data quantity for CPT and SFT stages during medical knowledge injection.
@@
Model CPT SFT English Chinese Japanese French Russian Spanish Average
InternLM-7B - 45K 44.07 64.62 37.19 24.92 58.20 44.97 45.67 +0.00
MMedLM-7B 25B 45K 49.88 70.49 46.23 36.66 72.27 54.52 55.01 +9.34
InternLM2-7B - 45K 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00
MMedLM2-7B 25B 45K 61.74 80.01 61.81 52.09 80.47 67.65 67.30 +8.71
Llama2-7B - 45K 43.36 50.29 25.13 20.90 66.80 47.10 42.26 +0.00
+Ours 76M 78K 49.41 65.15 36.68 35.21 69.14 50.62 51.04 +8.78
Llama2-13B - 45K 51.37 57.97 32.66 25.08 69.92 52.99 48.33 +0.00
+Ours 76M 78K 53.02 68.30 37.78 41.71 70.70 55.51 54.50 +6.17
InternLM2-7B - 45K 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00
+Ours 76M 78K 60.80 79.19 50.75 45.34 75.39 66.85 63.05 +4.46
@@
Ablation Studies. We conduct a comprehensive ablation study with the English split of the MMedBench dataset to evaluate our efficacy. Specifically, we take Llama2-7B as the backbone model, select the English textbooks [21] (with 26M tokens) for vanilla and our structure-aware continual pre-training, and use different QA samples for supervised fine-tuning. In particular, in Tab. 4, SFT refers to vanilla SFT with 10K English QA samples provided in MMedBench’s [35] training split, SSFT indicates structure-aware SFT on 10K training data, where the questions are the same as SFT while the answers are enhanced with knowledge structure explanation by Llama3-70B, as described in Sec. 3.3. SSFT* includes another 8K structure-aware QA syntheses by Llama3-70B, consisting of 18K entries for training. The training hyper-parameters follow the main experiment.
Results in Tab. 4 imply the vanilla CPT+SFT injection paradigm does bring considerable improvement of 1.73% in the English split, while our SCPT technique with vanilla SFT strategy presents a slightly higher accuracy of 46.50%. However, when combining SSFT with SCPT, our method immediately leads to a significant boost in the English split (49.96% v.s. 44.54%), demonstrating the efficacy and necessity of eliciting the learned structured knowledge to solve practical problems. Moreover, the supplementation of 8K extra QA syntheses surprisingly enhances the model performance on the other five subsets, resulting in a valuable gain on the average accuracy of 38.27%. This result demonstrates the knowledge transferability across different languages [24, 34]. After training with our SSFT strategy, LLMs can actively utilize the knowledge injected in one language to solve the problem in another language, further evidencing our superiority against the traditional SFT technique.
In addition, we observe that the commonly used RAG [25] strategy does not bring significant advantages to the MMedBench evaluation. The main reason lies in the gap between the pre-training corpus (comprising official knowledge statements from textbooks) and evaluated QA samples (originating from practical diagnosis records). To teach LLMs to apply relevant medical knowledge for reliable diagnosis answers, knowledge injection by CPT and SFT shows more advantages in this situation.
@@
Table 4: Ablation studies with Llama2-7B on the English subset of MMedBench. The best results are marked in bold, and the secondary results are marked with underlines.

### v2

@@ 4 Experiments
We design a comprehensive evaluation of our StructTuning through several experiments on two benchmarks. First, we investigate the free-form question-answering task on the LongBench (Bai et al., 2023) dataset, in order to verify the memorization and understanding of injected knowledge (the answer can be directly found in training corpora). Then, we delve into the multi-choice question-answering task on MMedBench (Qiu et al., 2024), to explore how LLMs apply the injected knowledge in basic medicine to determine the real-world diagnosis for patients with logical reasoning.
@@ 4.1 Preliminary Investigation on Free-form Question-Answering
Table 1: Recall evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench (Bai et al., 2023) dataset. The best results are marked in bold, and the secondary results are marked with underlines. The backbone model is Llama2-7B (Touvron et al., 2023b).
@@
OBQA - 39.7 44.3 17.5 30.1 35.5 11.9 9.6 26.9
CBQA CPT+SFT 20.7 35.3 20.6 29.9 32.1 18.9 12.0 24.2
SCPT+SFT 18.8 42.5 17.7 35.7 36.4 20.5 15.3 26.7
@@
Datasets and Tasks. LongBench (Bai et al., 2023) is a multi-task benchmark tailored for open-book reading comprehension evaluation, where LLMs generate answers to given questions based on one or several input passages. To focus on knowledge injection, we turn the open-book evaluation into a closed-book QA task, where LLMs are trained on contextual passages and queried with questions only. We choose 7 subsets with 1,350 test examples from LongBench for single- and multi-document question-answering evaluation, and the remaining synthetic or code-orientated tasks are eliminated. More details are described in Sec. A.1.
Evaluation Metrics. To quantify the knowledge memorization degree, we mainly report the recall (Zhu & Li, 2023b) for models’ outputs against ground-truth answers for the free-form QA tasks. In Sec. B.3 we also evaluate our method with the F1-score measure for a thorough comparison.
Investigated Models. For LongBench, we mainly investigate the knowledge injection to Llama2-7B (Touvron et al., 2023b) to compare the open- and closed-book QA performance.
Implementation Details. For closed-book QA, the Llama2-7B model is continually pre-trained on 10,476 passages for 3 epochs using a batch size of 128. The initial learning rate is 2e-5, decayed to 0 at the end of training with a cosine scheduler. Then, we query Llama3-70B to generate 2,700 SSFT QA examples and remove those with over 0.5 F1-Score similarity to test samples to prevent data leakage. We train the knowledge-injected models for 1 epoch to avoid overfitting.
Figure 6: An example for structure-aware responses.
Main Results. First, we use a CoT instruction of “Answer the question and explain why” to set a baseline for Open-Book QA (OBQA). As shown in Tab. 1, it presents a moderate knowledge comprehending capability in SingleDoc QA tasks (e.g., over 40% recall on Qasper and MFQA), and relatively poor knowledge coordination in MultiDoc QA subsets (e.g., around 10% recall on Musiq and Duzh) due to the attention drift on longer inputs.
Then, we try to inject passage content into LLMs with a conventional CPT+SFT paradigm to benchmark the Closed-Book QA (CBQA) baseline. Note that we use two types of question templates (refer to Sec. 3.3) during training, and use the CoT version during testing to elicit models’ memorized knowledge in their responses. However, Tab. 1 indicates such a knowledge injection approach is ineffective, as the knowledge recall is even lower than the OBQA baseline (i.e., 24.2% v.s. 26.9%).
On the contrary, our SCPT strategy has already achieved a considerable knowledge recall of 26.7%, which is higher than the CPT-SFT paradigm of 24.2% and approaching the OBQA baseline of 26.9%. It implies our structure-aware continually pre-trained model has successfully associated the relevant passages with their entire knowledge structure for the given question, which provides the knowledge path to seek targeted information to derive the answer.
Furthermore, our SSFT technique explicitly teaches the model to recall the learned knowledge and answer the questions, which further improves the knowledge recall to 31.5% and even surpasses the Open-Book QA setting. The results indicate the vanilla SFT strategy can only regularize LLMs’ response styles, while our SSFT could teach LLMs to utilize their knowledge (injected in the SCPT stage) to answer corresponding questions, as exemplified in Fig. 6.
Particularly, our method receives significant enhancements on MultiDoc QA subsets, which implies even though the total text content may exceed the models’ attention window and influence the OpenBook-QA, we can still chunk the content into several pieces and successfully inject the knowledge points into LLMs meanwhile preserving the whole knowledge structure.
@@
In addition, we also use lexical ROUGE-L (Lin, 2004) and semantic BERTScore (Zhang et al., 2020) to quantify the memorization of injected knowledge structures, by comparing the mindmap in models’ responses (as Fig. 6 displays) with ground-truth answers. The results in Fig. 1 indicate a relatively good memorization of the injected knowledge mindmap, emphasizing the efficacy of our SCPT strategy.
@@ 4.2 In-depth Evaluation for Multi-choice QA Application
Datasets and Tasks. MMedBench (Qiu et al., 2024) is a multilingual medical multi-choice QA benchmark, with 45,048 QA pairs for adapting LLMs to medical experts and 8,518 for testing. We collect 76M textbook corpora from MedTextBooks (Jin et al., 2020) and MMedC (Qiu et al., 2024) for medical knowledge injection and use the training/test split from MMedBench for SFT/evaluation. We also curate another two sizes of training sets (with 30M and 132M tokens) to validate our method’s scalability. Detailed setup is in Sec. A.2.
@@
Evaluation Metrics. For multi-choice QA in MMedBench (Qiu et al., 2024), we follow the default setting to calculate the accuracy on six language subsets, as well as the averaged scores. Metrics are computed by lexical exact-matching on models’ responses, rather than maximum token probabilities.
@@
Investigated Models. For MMedBench, we extend the investigated LLMs across model scales and architectures including Llama2-7B/13B (Touvron et al., 2023b), InternLM2-7B (Zheng et al., 2024), and the recent Llama3-8B (Dubey et al., 2024). We also compare our knowledge-injected models with other popular domain-specified LLMs, such as MedAlpaca (Han et al., 2023), ChatDoctor (Yunxiang et al., 2023), PMC-LLaMA (Wu et al., 2024), and MMedLM (Qiu et al., 2024) models.
@@
Implementation Details. We first train LLMs for 3 epochs on medical textbooks with a batch size of 128. Then, we ask Llama3-70B to create structure-aware explanations for existing 45K QA samples in MMedBench’s training split and 33K extra entries by traversing extracted knowledge structures. Syntheses with overlapped options in the test set are removed. In the SFT phase, the learning rate is set as 1e-6 to avoid overfitting on such an amount of SFT samples, as suggested by Qiu et al. (2024).
@@
Main Results. In Tab. 2, we present the overall performance across a series of LLMs on the testing split of MMedBench. The results demonstrate the promising enhancement achieved by our StructTuning technique, which translates into consistent improvements on the advanced InternLM2-7B (Zheng et al., 2024) and Llama3-8B (Dubey et al., 2024) models, and largely outperforms the previous domain-specific LLMs like PMC-LLaMA (Wu et al., 2024) and MedAlpaca (Han et al., 2023). Notably, our structure-aware knowledge injection approach, using merely 76M tokens curated from medical textbooks, achieves over 50% performance (4.46% v.s. 8.71%, 2.57% v.s. 4.96%) against the state-of-the-art MMedLM (Qiu et al., 2024) method, which is trained on a huge MMedC (Qiu et al., 2024) corpora of 25.5B tokens. Our approach shows much more efficiency in transforming pre-trained LLMs into domain experts by structured-aware knowledge injection, e.g., we only use 0.3% corpus to achieve comparable or even slightly better performance on English and Spanish subsets on top of Llama3-8B.
However, our method is currently unable to achieve 100% of knowledge injection efficacy against MMedLM, which may comprise two major reasons: (1) some knowledge in the training corpus has already been learned by the foundation model (e.g., a slight decrease in the Chinese subset), and (2) knowledge in the curated 76M training data cannot cover all testing scenarios (e.g., a considerable gap on the Japanese subset due to the absence of Japanese corpus, see Tab. A2). Therefore, we make a step to evaluate our approach’s scalability on various corpora sizes in Fig. 8.
@@
Table 2: Multiple-choice accuracy evaluation on MMedBench (Qiu et al., 2024). We report each model’s accuracy across six languages separately, with “Average” denoting the mean score over six languages. We also compare the data requirement (“#Token”) for medical knowledge injection.
@@
Model English Chinese Japanese French Russian Spanish Average #Token
ChatDoctor 43.52 43.26 25.63 18.81 62.50 43.44 39.53 -
PMC-LLaMA 47.53 42.44 24.12 20.74 62.11 43.29 40.04 -
MedAlpaca 46.74 44.80 29.64 21.06 59.38 45.00 41.11 -
Llama2-7B 43.36 50.29 25.13 20.90 66.80 47.10 42.26 -
InternLM-7B 44.07 64.62 37.19 24.92 58.20 44.97 45.67 -
InternLM2-7B 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00 -
+MMed 61.74 80.01 61.81 52.09 80.47 67.65 67.30 +8.71 25.5B
+Ours 60.80 79.19 50.75 45.34 75.39 66.85 63.05 +4.46 76M
Llama3-8B 63.86 78.23 48.24 50.80 71.48 64.15 62.79 +0.00 -
+MMed 66.06 79.25 61.81 55.63 75.39 68.38 67.75 +4.96 25.5B
+Ours 66.77 77.44 53.27 51.61 74.61 68.49 65.36 +2.57 76M
Figure 8: Comparison of knowledge injection approaches.
Approach’s Scalability. In Tab. 2, we use roughly 0.3% of the 25.5B tokens in MMedC Qiu et al. (2024) to evaluate knowledge injection methods. Here, another two training corpora sizes are curated: 30M and 132M, which take around 0.1% and 0.5% of 25.5B tokens respectively. We compare the vanilla CPT-SFT paradigm and our SCPT-SSFT strategy across those data settings The backbone LLM is InternLM2-7B (Zheng et al., 2024), and the training settings follow the main experiments. According to Fig. 8, our method consistently surpasses the vanilla paradigm by a large margin, emphasizing the efficacy and efficiency of domain knowledge injection. In particular, we can fit two performance-ratio scaling curves from the data points in Fig. 8 as:
$p_{v}\approx-0.04(\log r)^{2}+13.3\log r+100.0;\quad p_{s}\approx-1.11(\log r)^{2}+7.63\log r+133.0$ (3)
where $p_{v}$ and $p_{s}$ denote the relative performance enhancement (%) for vanilla and structure-aware knowledge injection, and $r$ refers to the corpus ratio. The scaling law indicates we can achieve comparable performance (100%) with only 5% of the total training corpus, significantly reducing the costs for LLMs’ domain adaptation. On the other hand, it also indicates our method may lead to 133% enhancement with a further 100% comprehensive data utilization. However, we have not empirically verified those predictions due to time and resource constraints, which we view as a temporal limitation of our current work and leave it to future investigations.
Approach’s Generalization. In addition to the InternLM2 and Llama3 models in previous experiments, we also investigate the generalization ability of our SCPT+SSFT paradigm on Llama2 (Touvron et al., 2023b) model series. As shown in Tab. 3, our method leads to consistently significant improvements on 7B (+8.78%) and 13B (+6.17%) backbone models. The results further demonstrate the generalizability and scalability of our StructTuning strategy across model architectures and sizes.
Table 3: Structure-aware knowledge injection to Llama2 (Touvron et al., 2023b) model series.
Model English Chinese Japanese French Russian Spanish Average
Llama2-7B 43.36 50.29 25.13 20.90 66.80 47.10 42.26
+Ours 49.41 65.15 36.68 35.21 69.14 50.62 51.04
Llama2-13B 51.37 57.97 32.66 25.08 69.92 52.99 48.33
+Ours 53.02 68.30 37.78 41.71 70.70 55.51 54.50
@@
Ablation Studies. To further demonstrate the efficacy of each component of our StructTuning method, we conduct a comprehensive ablation study with the English split of the MMedBench dataset. Specifically, we take Llama2-7B as the backbone model, select the English textbooks (Jin et al., 2020) (with 26M tokens) for vanilla and our structure-aware continual pre-training, and use different QA samples for supervised fine-tuning. In Tab. 4, “SFT” refers to vanilla SFT with 10K English QA samples provided in MMedBench’s (Qiu et al., 2024) training split, “SSFT” indicates structure-aware SFT on 10K training data, where the questions are the same as “SFT” while the answers are enhanced with knowledge structure explanation by Llama3-70B, as described in Sec. 3.3. “SSFT*” includes another 8K structure-aware QA syntheses by Llama3-70B, consisting of 18K entries for training. The training hyper-parameters follow the main experiment.
@@
Table 4: Ablation studies with Llama2-7B on the English subset of MMedBench.
@@
In Tab. 4, vanilla CPT+SFT injection paradigm does bring an improvement of 1.73% in the English split, while our SCPT technique with vanilla SFT strategy presents a slightly higher accuracy of 46.50%. However, when combining SSFT with SCPT, our method immediately leads to a significant boost in the English split (49.96% v.s. 44.54%), demonstrating the efficacy and necessity of eliciting the learned structured knowledge to solve practical problems. Moreover, the supplementation of 8K extra QA syntheses (“SSFT*”) surprisingly enhances the model performance on the other five subsets, resulting in a valuable gain on the average accuracy to 38.27%. This result demonstrates the knowledge transferability across different languages (Lai et al., 2023; Qin et al., 2024). After training with SSFT, LLMs can actively utilize the knowledge injected in one language to solve the problem in another language, evidencing our superiority against the traditional SFT technique. In Sec. B.4, we follow Liu et al. (2024b) to randomly generate another 8K SFT pairs for further comparison, where our structure-aware SFT syntheses are verified to better enhance LLMs’ knowledge application.
In addition, we observe that the commonly used RAG (Lewis et al., 2020) strategy does not bring significant advantages to the MMedBench evaluation. The main reason lies in the gap between the pre-training corpus (comprising official knowledge statements from textbooks) and evaluated QA samples (originating from practical diagnosis records). Knowledge injection by (S)CPT and (S)SFT shows more advantages in this situation.